OpenStudy (anonymous):

find the derivative : f(x)=-5lnx

jimthompson5910 (jim_thompson5910):

the derivative of ln(x) is 1/x and this is something you should memorize or write on a cheatsheet (or notecard)

jimthompson5910 (jim_thompson5910):

you pull out the constant, so you ignore that for now and derive ln(x) to get 1/x then you reintroduce the constant back in

OpenStudy (anonymous):

alright i've seen that I just dont understand how it plays in with the -5

jimthompson5910 (jim_thompson5910):

that's a constant that can be ignored when you derive

jimthompson5910 (jim_thompson5910):

any constants can be pulled out

jimthompson5910 (jim_thompson5910):

so you would derive y = 22*ln(x), y = 18*ln(x) y = -33*ln(x) all the same way

OpenStudy (anonymous):

ooooooohhh okay

OpenStudy (anonymous):

-5*1/x

jimthompson5910 (jim_thompson5910):

which becomes -5/x

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

It'd be nice if my text would've just said that instead of going around in circles

jimthompson5910 (jim_thompson5910):

it probably did somewhere, but it just gave a convoluted way of saying it (with lots of notation attached)

OpenStudy (mathstudent55):

If k = constant and U = expression is x, then d/dx(kU) = k * d/dx(U) d/dx(-5lnx) = -5* d/dx(lnx) = -5 * (1/x) = -5/x